OpenStudy (anonymous):

If 1/x > 2 and 1/y > 3, prove that 2x + y > 8xy. x,y - belong to R*.

OpenStudy (anonymous):

 and I tried to get the inequality from the hypothesis.

OpenStudy (anonymous):

\[\frac{ 1 }{ x }>2 \rightarrow x >\frac{ 1 }{ 2 }\]\[\frac{ 1 }{ y }>3 \rightarrow y >\frac{ 1 }{ 3 }\]Plug in to the \[2x+y >8xy\](I dunno if this is right)

OpenStudy (anonymous):

shouldn't there be equality between x and 1/2 in order to plug in to the inequality?

OpenStudy (anonymous):

ohhh yeah, ehehehe